A tuning fork of frequency 480 Hz resonates with a tube closed at one end of length 16 cm and diameter 5 cm in fundamental mode. Calculate velocity of sound in air.    

Grade:10

2 Answers

Jitender Pal
9 years ago
Hello Student,
Please find the answer to your question
KEY CONCEPT : In the fundamental mode
(ℓ + 0.6r) = ℷ/4 = v/4f ⇒ v = 4f (ℓ + 0.6r) = 336 m/s.

Thanks
Jitender Pal
askIITians Faculty
ankit singh
3 years ago
As tension increases, frequency increases in the string as fT
where T is tension,
 
increase in tension in the string frequency of the string increases and as given in problem ,
beats decreases so frequency of string must be less than frequency of fork.
 
So, Beat =frequency of fork frequency of string
 
10=480 Hzfs
 
frequency of string = 470 Hz